Alvin paddled for 4 hourswith a 3km/h current to reach a campsite. The return trip against the same current took 7 hours. What is the speed of the boat in still water? What is the speed of the boat against the current?
Answer by swincher4391(1107)

Let the boat speed in still water be x km/hr.
We know the current is 3 km/h so:
4(x+3) = 7(x-3)
4x + 12 = 7x - 21
-3x = -33
x = 11 km/h